Last week we passed a significant milestone at the HET. We brought the tracker bridge out from Austin and lifted it into place. It went very smoothly and you can find a video of the lift at http://het.as.utexas.edu/wfu/.
This week we are cabling up the last of the X and Y cables. Then we will test the X motion of the tracker including the X-skew prevention under full closed loop control. Late in the week and extending into the next expect to start the installation of the Y drive system (including the so called anti-gravity drive). We hope to be able to gain back some of the two weeks of time we have slipped on the schedule by also putting the hex drive together at the same time.
Last week week we finished clearing out the optics tables and floor for the MRS decommissioning. We also met with the HZPF team to discuss their preliminary plan for their new enclosure in the spectrograph room. Much of the staff were preoccupied with finishing preparation for the tracker lift. The crane arrived late on Friday afternoon (3 days early).
This week the Crane Services will setup for their work on Monday and will take out the lower and upper X rails that were used for shipping. We are now done with them (since the lower and upper X components are assembled and working on the telescope). On Tuesday the tracker should arrive from Austin in the morning and if the weather cooperates Crane Services will lift it into place. We expect that this will take much of the day. We hope it does not take longer because the winds are predicted to rise on Wednesday and peak at “Destructive Winds” on Thursday. Once the tracker bridge is in place we will begin cabling it and start testing moves over the next weeks.
Have a look at the link below for the promised video of the Last week’s work.
Cable Wrap Installation
Notice the use of hard hats in the dome (good safety practice) but it is harder to notice that everyone is freezing cold.
Last week we finished the cabling of the DAS box and populating most of the cables in the Azimuth wrap. We pulled in Resident Astronomer’s, Telescope Operators, the full day staff and even some people from McDonald Observing support. Thanks to everyone for the hours of hard work in the freezing temperatures. With so many people working so hard we neglected to take any photos of the work but I think we may try to put something together using the dome safety cameras.
This week we are getting ready for the new tracker arrival next week. This means we need to get the upper and lower X in full working order with limit switches in place, using the encoders for active feedback and have our emergency-stop system integrated into the new Tracker hardware. In addition to this work, we will be finishing decommissioning of the MRS enclosure to make room for the first group of VIRUS spectrographs to be shipped to West Texas sometime in the next month. We will have several different contractors on site this week. One connecting our large UPS system for the tracker, and another mounting the cryogenic lines from the site of the huge liquid nitrogen tank to the inner pier.
Last week we were successful in moving and aligning both the lower and upper X-slew axes. We also began the really massive task of cabling the tracker. Based on the rates of cabling we will likely be cabling until Thursday afternoon this week. This requires a team of ~10 people to connect the cable up at the DAS box and then down the cable trays, through the cable wrap, down to the spectrograph room and then through the spectrograph room cable trays into the lower electrical room.
Once the cabling is done we will start testing full motion of both X axes including limit switches and simultaneous moves of upper and lower x. All of this in preparation for the new tracker lift on the 18th (weather permitting).